OpenStudy (anonymous):

So ya, this question is in relation to applications of derivatives, I'm just not sure how to solve. Any help would be awesome~ A balloon is rising vertically above a level, straight road at a constant rate of 1 ft/sec. Just when the balloon is 130ft above the ground, a bicycle moving at a constant rate of 17ft/sec passes under it. How fast is the distance s(t) between the bicycle and balloon increasing 6 sec later?

OpenStudy (anonymous):

dh/dt = 1 dx/dt=17

OpenStudy (anonymous):

h is height x represent position of bike, so dx reprent speed

OpenStudy (anonymous):

now pythagorean triangle x^2 + h^2 = d^2 \[d=\sqrt{h^2 + x^2}\]

OpenStudy (anonymous):

Take partial from here

OpenStudy (anonymous):

haven't done this for a while... so I probably made a mistake or two (or 3). but here's my attempt... if we differentiate the relation (pythagorean theorem above) we get 2(dh/dt)h+2(dx/dt)x=2d(dd/dt) and just sub in the values evaluated at t=6. h(6)=136, x(6)=17*6 then solve for dd/dt.

OpenStudy (anonymous):

Yep, It looks right

OpenStudy (anonymous):

yeah, it is better idea to take partial of d^2 =h^2 + x^2 gj

OpenStudy (anonymous):

alright, i see how it works ands got an answer (sr about no reply earlier, net cut off -_-)... alright, thanks a mil u 2.